question-A man has some hens and cows.If the number of heads be 48 and the number of feet equals 140,then the number of hens will be?
Answer by mananth(16946)

cows x numbers
hens y numbers

x+y=48 .............1
4x+2y=140 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 1
-2x-2y=-96
4x+2y= 140
Add the two equations
2x=44
/2
x= 22
plug value of x in (1)
x+y=48
22+y=48
y=48-22
y = 26

cows 22 numbers
hens 26 numbers
m.ananth@hotmail.ca